OpenStudy (anonymous):

Use the arc length formula and the given information to find r. s = 12 cm, θ = 36°; r = ?

OpenStudy (anonymous):

30/pi or 60/pi im not sure please help wich 1 i think its deff 30/pi

OpenStudy (anonymous):

the formula is s = r*theta, where theta is given in RADIANS. so you should first convert that degree measure into radians then you can use the formula.

OpenStudy (anonymous):

check

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i did this

OpenStudy (anonymous):

\(\large \theta = 36 \cancel {deg} \cdot \frac{\pi rad}{180 \cancel {deg}}=\frac{ \pi }{5} \text radians \)

OpenStudy (anonymous):

so now, \(\large 12=\frac{\pi}{5}\cdot r \rightarrow r=\frac{60}{\pi}\)

OpenStudy (anonymous):

oh i see thank you so much